write an equation of the circle described: center is (2,3); passes through (5,6).
Answer by venugopalramana(3286)

write an equation of the circle described: center is (2,3); passes through (5,6).
EQN OF CIRCLE WITH CENTRE AS (H,K)IS
(X-H)^2+(Y-K)^2=R^2
(X-2)^2+(Y-3)^2=R^2...SINCE THIS GOES THROUGH (5,6)..WE HAVE
(5-2)^2+(6-3)^2=R^2=3^2+3^2=18
HENCE EQN.OF CIRCLE IS
(X-2)^2+(Y-3)^2=18